OpenStudy (anonymous):

I am trying to solve a geometry word problem; can someone help me. "The perimeter of an equilateral triangle is 7 centimeters more than the perimeter of a square, and the side of the triangle is 5 centimeters longer than the side of a square. Find the side length of the triangle. (note: an equilateral triangle has three equal side lengths)

OpenStudy (anonymous):

let the side of square be = x cm so its perimeter = 4x cm side of equilateral triangle = (x + 5) cm so its perimeter = 3*(x + 5) cm = (3x + 15) cm ATQ 3x + 15 = 4x + 7 15 - 7 = 4x - 3x 8 = x Hence, length of side of triangle = x + 5 = 8 + 5 = 13 cm

OpenStudy (anonymous):

hey eball, i need one more fan to complete the number of 100.... can u become my fan pls....☺

OpenStudy (anonymous):

Many THANKS !!!!!! :))

OpenStudy (anonymous):

THANK YOU! :-) YOU'RE AWESOME